# Value-Guided Search for Efficient Chain-of-Thought Reasoning

### v1

@@ Abstract
In this paper, we propose a simple and efficient method for value model training on long-context reasoning traces. Compared to existing process reward models (PRMs), our method does not require a fine-grained notion of “step,” which is difficult to define for long-context reasoning models. By collecting a dataset of 2.5 million reasoning traces, we train a 1.5B token-level value model and apply it to DeepSeek models for improved performance with test-time compute scaling. We find that block-wise value-guided search (VGS) with a final weighted majority vote achieves better test-time scaling than standard methods such as majority voting or best-of-$n$. With an inference budget of 64 generations, VGS with DeepSeek-R1-Distill-1.5B achieves an average accuracy of $45.7\%$ across four competition math benchmarks (AIME 2024 & 2025, HMMT Feb 2024 & 2025), reaching parity with o3-mini-medium. Moreover, VGS significantly reduces the inference FLOPs required to achieve the same performance of majority voting. Our dataset, model and codebase are open-sourced at https://github.com/kaiwenw/value-guided-search.
@@
Figure 1: Performance and Efficiency of Value Guidance: (Left) Value-guided search improves the overall quality of DeepSeek-R1-Distill responses across four combined competition math benchmarks (AIME & HMMT). The inference budget for 1.5B, 7B and 14B are $256$, $128$ and $64$ generations, respectively. (Right) Value-guided search also reduces the inference FLOPs required to achieve the same accuracy levels as majority voting, a standard TTC scaling baseline, showing value-guidance is promising for improving efficiency.
@@ 1 Introduction
We propose value-guided search (VGS) – a block-level search method guided by a token-level value model – as a promising approach to scale TTC for reasoning models. In Section 2, we present an effective pipeline for value model training on tasks with outcome labels, such as competition math. Our data pipeline collects solution prefixes from various models and then, starting from random prefixes, generates completed solutions using a lean reasoning model (e.g., DeepSeek-R1-Distill-1.5B). Notably, our data collection does not require a pre-defined notion of step and is more efficient than existing techniques [44, 27]. With this pipeline, we collect a dataset of 2.5 million math reasoning traces (over $30$ billion tokens) from a filtered subset of the OpenR1-Math dataset [2]. Then, we train a 1.5B token-level value model called DeepSeek-VM-1.5B by regressing (via classification) the final reward of the completed solution.
Next, in Section 3, we apply our value model to perform block-wise search with DeepSeek models [18] on competition math, where we evaluate on four prestigious high school math competitions in the US (AIME 2024 & 2025 and HMMT 2024 & 2025). Our experiments show that block-wise VGS significantly improves TTC compared to majority voting or weighted majority voting, strong baselines from the literature [45, 6]. We also show that VGS with DeepSeek-VM-1.5B leads to higher performance than searching with state-of-the-art PRMs, demonstrating that our value model can provide better feedback. When given an inference budget of $64$ generations, VGS on DeepSeek-R1-Distill-Qwen-1.5B can outperform o3-mini-medium, and VGS on DeepSeek-R1-Distill-Qwen-14B (total size with value model is 15.5B) is on par with DeepSeek-R1 (671B) on our competition math evaluations (Fig. 1 left). Moreover, we show that VGS reduces the amount of inference compute required to attain the same performance as majority voting (Fig. 1 right). In summary, we find that block-wise VGS not only improves the performance ceiling of reasoning models, but also reduces the amount of inference compute required to match the performance of standard TTC methods. Our contributions are summarized below:
@@
    A simple recipe for token-level value model training that does not require a pre-defined notion of “step” and scales to long-context reasoning traces (Section 2).
@@
    Block-wise search, guided by our 1.5B value model, achieves impressive performance on four challenging math competitions, outperforming standard TTC methods (e.g., best-of-$N$, majority voting) and search with existing PRMs (Section 3).
@@
Please see Appendix A for a detailed discussion of related works.
@@ 2 Methods
We present an end-to-end pipeline for training a token-level value model and applying it to guide block-wise search. In Section 2.1, we introduce necessary notation and present a regression-via-classification algorithm for learning the token-level value model [14]. Then, in Section 2.2, we outline an efficient data pipeline for creating our dataset of 2.5 million reasoning traces from DeepSeek models. Finally, in Section 2.3, we describe several TTC methods and baselines, e.g., best-of-$N$, (weighted) majority voting and search algorithms that can leverage our value model. While we focus on competition math in this paper, we remark that our pipeline can in principle be applied to any task with automated outcome supervision (e.g., a reward model). In Appendix B, we summarize a simple recipe for applying VGS to other such domains.
Figure 2: Summary of Methods. (Left) Diagrams how we collect multiple roll-ins (grey circles representing tokens) per problem, and branch off multiple roll-outs per roll-in at random points. The class label for each roll-out token is the outcome label at the very end. (Right) Shows the beam search process (beam width $2$ and budget $4$) guided by a value model.
@@ 2.1 Learning Algorithm for Value Model
We describe our training process for a language value model by performing regression via classification [7]. Let $\mathcal{V}$ be the vocabulary and let $\mathcal{S}=\bigcup_{n\in\mathbb{N}}\mathcal{V}^{n}$ denote the input sequence space. Given a problem prompt $x\in\mathcal{S}$ and a response $y\in\mathcal{S}$, let $\kappa=\Gamma(x,y)\in[0,1,\dots,K-1]$ denote its class label, where $K$ is the number of classes. Furthermore, let $r=R(x,y)$ denote the scalar reward, which we assume to be binary since we focus on competition math (see Appendix B for the general case). For our value model, $\kappa=2$ if the response is an incomplete generation (i.e., exceeds max generation length), $\kappa=0$ if the response finished and is incorrect, and $\kappa=1$ if the response finished and is correct. Thus, the event that $\kappa=1$ corresponds to $r=1$ (correct answer), and $\kappa\in\{0,2\}$ corresponds to $r=0$ (incorrect or exceeds max length). We adopt this convention in the rest of the paper. We remark that regression-via-classification is a standard approach that leads to better down-stream decision making than regressing via squared error [7, 20, 17, 3, 42, 43].
@@ 2.2 Dataset Creation Process
Post-Filtering. We filter out problems where all $56$ roll-outs for that problem were incomplete or incorrect (i.e., has reward 00). This post-filtering removes any ambiguous or unanswerable problems that we missed during pre-filtering, and also removes problems that are too difficult for $\pi^{\text{ref}}$ and do not provide a useful learning signal. This step filters roughly $10\%$ of problems, yielding a final dataset of 2.5 million datapoints.
Notably, our approach does not require a fine-grained notion of step and our data collection is cheaper than existing PRM techniques. Specifically, Lightman et al. [23] used per-step annotations by human experts, Zhang et al. [50] used per-step annotations via LLM-as-a-Judge, and Wang et al. [44] used multiple Monte Carlo roll-outs at every step. Since the number of newlines in reasoning CoT traces can grow very quickly, per-step labels are quite expensive to collect for reasoning models. In contrast, our approach only requires a handful of roll-ins (from any policy) and roll-outs (from $\pi^{\text{ref}}$) per problem, and this number can be flexibly tuned up or down to trade-off data coverage and data collection cost. Please refer to Appendix D for further details on each step. We also release our filtering code and datasets to support future research.
@@ 2.3 Algorithms for Test-Time Compute and Search
Beam Search. One weakness of BFS is that parallel blocks are correlated because they share the same prefix, which limits diversity. Beam search with width $w$ (Algorithm 1) is a generalization that keeps $B=N/w$ (assume to be integer) partial responses and branches $w$ parallel blocks from each one [26, 5, 40, 6, 36]. Given a prompt $x$, beam search first generates $N$ parallel blocks. However, unlike BFS, beam search keeps the top $B$ beams with the highest scores, and then samples $w$ parallel blocks per beam at the next step. Since $B\times w=N$ blocks are sampled at each step, the compute budget is also $N$. We illustrate beam search with $N=4$ and $w=2$ in Fig. 2 (right).
@@
We remark a crucial detail of beam search and DVTS is how the final set of beams/responses are aggregated. Prior works [6, 36, 34] select the response with the highest score, which is analogous to a final best-of-$N$ (BoN; Algorithm 2). Instead, we found that taking a weighted majority vote (WMV; Algorithm 3) led to much better performance, which is demonstrated by Fig. 3 (left).
Computational Efficiency of Block-wise Search. Since value scores are only used at the end of each block or the end of the whole response, the FLOPs required for block-wise value model guidance is a tiny fraction ($\ll 1\%$) of the generation cost from $\pi$. We compute FLOP estimates in Appendix H to concretely show this.
@@ 3 Experiments
We extensively evaluate value-guided search (VGS) with our 1.5B value model DeepSeek-VM-1.5B, focusing on guiding the CoT reasoning of DeepSeek models [18]. The best VGS setup for our value model is beam search with final WMV aggregation, beam width 2, block size 4096 and with DVTS (for larger inference budgets). We show this setup outperforms other test-time compute methods (e.g., MV, WMV, BoN) and other scoring models (e.g., existing 7B PRMs and a 1.5B Bradley-Terry reward model trained on our dataset). We remark our search results use a fixed beam width and block size for all problems; this is more practical than prior works on “compute-optimal scaling” which vary search parameters for each problem and require estimating each problem’s difficulty [6, 36, 25]. Please see Appendices E and F for additional details on value model training and inference.
Benchmarks. We evaluate on the 2024 and 2025 editions of the American Invitational Mathematics Exam (AIME) and the February Harvard-MIT Mathematics Tournament (HMMT). [^2] https://maa.org/maa-invitational-competitions and https://www.hmmt.org Both AIME and HMMT are prestigious high school math competitions in the US that have also been used to evaluate frontier LLMs [31, 18, 1]. We use AIME I & II and the individual part of HMMT, yielding $30$ problems per competition. To mitigate overfitting on a single widely used benchmark, we report the overall averaged accuracy unless otherwise stated. Per-benchmark plots are relegated to Appendix C.
Baseline Models. We evaluate two state-of-the-art 7B PRMs with distinct training styles: Math-Shepherd-Mistral-7B-PRM [44] and Qwen2.5-Math-PRM-7B [50]. Math-Shepherd uses Monte-Carlo roll-outs from each step to estimate per-step value while the Qwen2.5 PRM uses LLM-Judge annotation for each step, similar to the per-step human annotation of PRM800K [23]. As a step-level value model, Math-Shepherd-PRM-7B is more related to our token-level value model. Finally, we also evaluate a 1.5B Bradley-Terry (BT) [8] model, called DeepSeek-BT-1.5B, which we trained using our dataset (see Appendix G for training details).
MV@256 $71.0\pm 3.5$ $38.9\pm 1.9$ $24.4\pm 1.7$ $24.3\pm 2.9$ $39.7\pm 1.2$
Test-time scaling larger models with our DeepSeek-VM-1.5B
MV w/ DeepSeek-7B ($N=128$) $77.1\pm 1.1$ $56.5\pm 1.6$ $34.7\pm 1.6$ $33.8\pm 2.5$ $50.5\pm 0.9$
@@
DeepSeek-1.5B Pass@1 $28.2\pm 6.1$ $22.4\pm 4.1$ $13.9\pm 4.2$ $13.0\pm 3.9$ $19.4\pm 1.1$
Table 1: (Top) Weighted majority vote (WMV) and VGS results for DeepSeek-1.5B with an inference budget of $N=256$, using various scoring models. (Middle) Compares MV and VGS for larger DeepSeek models guided with our DeepSeek-VM-1.5B. (Bottom) Lists performance of DeepSeek models and strong close-sourced reasoning models. For VGS, $\pm$ indicates standard deviation across 3 seeds; for MV, WMV, Pass@N, $\pm$ denotes bootstrap with $100$ repetitions. We bold the highest avg. accuracy and underline second highest. Section C.1 contains more baselines.
3.1 Main Results (Table 1)
In the top section of Table 1, we fix the generator to DeepSeek-1.5B [^3] Throughout the paper, we use DeepSeek-XB as shorthand for DeepSeek-R1-Distill-Qwen-XB. and test-time budget to $N=256$, and compare VGS to weighted majority voting (WMV), using our value model, the BT model and baseline PRMs. We see that VGS and WMV with DeepSeek-VM-1.5B achieve the two highest scores, outperforming the BT reward model and prior PRMs. This shows that our value model is not only a strong outcome reward model (ORM) but also an effective value model for guiding search. Notably, with a budget of $256$, our 1.5B value model can guide DeepSeek-1.5B (total parameter count is 3B) to reach parity with the pass@1 of o3-mini-medium, a strong math reasoning model from OpenAI. Intriguingly, while DeepSeek-BT-1.5B was only trained as an ORM, we find that VGS also improves performance relative to WMV, suggesting that BT models may also provide meaningful block-wise feedback to guide search. We also observe that accuracies for the 7B baseline PRMs (MathSheperd and Qwen2.5-Math) are only slightly higher than MV@256 and do not improve with search, which suggests that these PRMs are likely out-of-distribution (OOD) for the long CoTs generated by DeepSeek-1.5B.
In the middle section of Table 1, we guide stronger DeepSeek models with sizes 7B and 14B, and compare VGS to MV, a standard TTC method that does not use an external scoring model. We see that VGS again achieves higher accuracy than MV for both 7B and 14B, which suggests that DeepSeek-VM-1.5B is also useful in guiding the CoT of stronger DeepSeek models. However, we observe that the gap between VGS and MV becomes smaller for larger DeepSeek models, suggesting that DeepSeek-14B CoTs may be becoming OOD for our value model, which was trained on DeepSeek-1.5B CoTs. To guide more capable models, new value models should be trained on rollouts from similarly capable models; we however do not foresee this being a practical concern given the scalability of our training process (described in Section 2 and summarized in Appendix B). Finally, we note that the performance of all models on AIME-24 is consistently higher than other competitions, suggesting the importance of evaluating on diverse and newer competitions to reduce risk of overfitting or data contamination.

### v2

@@ Abstract
In this paper, we propose a simple and efficient method for value model training on long-context reasoning traces. Compared to existing process reward models (PRMs), our method does not require a fine-grained notion of “step,” which is difficult to define for long-context reasoning models. By collecting a dataset of 2.5 million reasoning traces, we train a 1.5B token-level value model and apply it to DeepSeek models for improved performance with test-time compute scaling. We find that block-wise value-guided search (VGS) with a final weighted majority vote achieves better test-time scaling than standard methods such as majority voting or best-of-$n$. Moreover, VGS significantly reduces the inference FLOPs required to achieve the same performance of majority voting. Our dataset, model and codebase are open-sourced at https://github.com/kaiwenw/value-guided-search.
@@
Figure 1: Performance and Efficiency of Value Guidance: (Left) Value-guided search improves the overall quality of DeepSeek-R1-Distill responses across combined competition math benchmarks (AIME 24, 25 & HMMT Feb 24, 25). The inference budget for 1.5B, 7B and 14B are $256$, $128$ and $64$ generations, respectively. (Right) Value-guided search also reduces the inference FLOPs required to achieve the same accuracy levels as majority voting, a standard TTC scaling baseline, showing value-guidance is promising for improving efficiency.
@@ 1 Introduction
We propose value-guided search (VGS) – a block-level search method guided by a token-level value model – as a promising approach to scale TTC for reasoning models. In Section˜2, we present an effective pipeline for value model training on tasks with outcome labels, such as competition math. Our data pipeline collects solution prefixes from various models and then, starting from random prefixes, generates completed solutions using a lean reasoning model (e.g., DeepSeek-R1-Distill-1.5B). Notably, our data collection does not require a pre-defined notion of step and is more efficient than existing techniques [45, 28]. With this pipeline, we collect a dataset of 2.5 million math reasoning traces (over $30$ billion tokens) from a filtered subset of the OpenR1-Math dataset [2]. Then, we train a 1.5B token-level value model called DeepSeek-VM-1.5B by regressing (via classification) the final reward of the completed solution.
Next, in Section˜3, we apply our value model to perform block-wise search with DeepSeek models [19] on competition math, where we evaluate on four prestigious high school math competitions in the US (AIME 2024 & 2025 and HMMT 2024 & 2025). Our experiments show that block-wise VGS significantly improves TTC compared to majority voting or weighted majority voting, strong baselines from the literature [46, 6]. We also show that VGS with DeepSeek-VM-1.5B leads to higher performance than searching with state-of-the-art PRMs, demonstrating that our value model can provide better feedback. When given an inference budget of $64$ generations, VGS on DeepSeek-R1-Distill-Qwen-14B (total size with value model is 15.5B) is on par with DeepSeek-R1 (671B) on our competition math evaluations (Fig.˜1 left). Moreover, we show that VGS reduces the amount of inference compute required to attain the same performance as majority voting (Fig.˜1 right). In summary, we find that block-wise VGS not only improves the performance ceiling of reasoning models, but also reduces the amount of inference compute required to match the performance of standard TTC methods. Our contributions are summarized below:
@@
    A simple recipe for token-level value model training that does not require a pre-defined notion of “step” and scales to long-context reasoning traces (Section˜2).
@@
    Block-wise search, guided by our 1.5B value model, achieves impressive performance on four challenging math competitions, outperforming standard TTC methods (e.g., best-of-$N$, majority voting) and search with existing PRMs (Section˜3).
@@
Please see Appendix˜A for a detailed discussion of related works.
@@ 2 Methods
We present an end-to-end pipeline for training a token-level value model and applying it to guide block-wise search. In Section˜2.1, we introduce necessary notation and present a regression-via-classification algorithm for learning the token-level value model [14]. Then, in Section˜2.2, we outline an efficient data pipeline for creating our dataset of 2.5 million reasoning traces from DeepSeek models. Finally, in Section˜2.3, we describe several TTC methods and baselines, e.g., best-of-$N$, (weighted) majority voting and search algorithms that can leverage our value model. While we focus on competition math in this paper, we remark that our pipeline can in principle be applied to any task with automated outcome supervision (e.g., a reward model). In Appendix˜B, we summarize a simple recipe for applying VGS to other such domains.
Figure 2: Summary of Methods. (Left) Diagrams how we collect multiple roll-ins (grey circles representing tokens) per problem, and branch off multiple roll-outs per roll-in at random points. The class label for each roll-out token is the outcome label at the very end. (Right) Shows the beam search process (beam width $2$ and budget 44) guided by a value model.
@@ 2.1 Learning Algorithm for Value Model
We describe our training process for a language value model by performing regression via classification [7]. Let $\mathcal{V}$ be the vocabulary and let $\mathcal{S}=\bigcup_{n\in\mathbb{N}}\mathcal{V}^{n}$ denote the input sequence space. Given a problem prompt $x\in\mathcal{S}$ and a response $y\in\mathcal{S}$, let $\kappa=\Gamma(x,y)\in[0,1,\dots,K-1]$ denote its class label, where $K$ is the number of classes. Furthermore, let $r=R(x,y)$ denote the scalar reward, which we assume to be binary since we focus on competition math (see Appendix˜B for the general case). For our value model, $\kappa=2$ if the response is an incomplete generation (i.e., exceeds max generation length), $\kappa=0$ if the response finished and is incorrect, and $\kappa=1$ if the response finished and is correct. Thus, the event that $\kappa=1$ corresponds to $r=1$ (correct answer), and $\kappa\in\{0,2\}$ corresponds to $r=0$ (incorrect or exceeds max length). We adopt this convention in the rest of the paper. We remark that regression-via-classification is a standard approach that leads to better down-stream decision making than regressing via squared error [7, 21, 17, 3, 43, 44].
@@ 2.2 Dataset Creation Process
Post-Filtering. We filter out problems where all $56$ roll-outs for that problem were incomplete or incorrect (i.e., has reward 0). This post-filtering removes any ambiguous or unanswerable problems that we missed during pre-filtering, and also removes problems that are too difficult for $\pi^{\text{ref}}$ and do not provide a useful learning signal. This step filters roughly $10\%$ of problems, yielding a final dataset of 2.5 million datapoints.
Notably, our approach does not require a fine-grained notion of step and our data collection is cheaper than existing PRM techniques. Specifically, Lightman et al. [24] used per-step annotations by human experts, Zhang et al. [51] used per-step annotations via LLM-as-a-Judge, and Wang et al. [45] used multiple Monte Carlo roll-outs at every step. Since the number of newlines in reasoning CoT traces can grow very quickly, per-step labels are quite expensive to collect for reasoning models. In contrast, our approach only requires a handful of roll-ins (from any policy) and roll-outs (from $\pi^{\text{ref}}$) per problem, and this number can be flexibly tuned up or down to trade-off data coverage and data collection cost. Please refer to Appendix˜D for further details on each step. We also release our filtering code and datasets to support future research.
@@ 2.3 Algorithms for Test-Time Compute and Search
Beam Search. One weakness of BFS is that parallel blocks are correlated because they share the same prefix, which limits diversity. Beam search with width $w$ (Algorithm˜1) is a generalization that keeps $B=N/w$ (assume to be integer) partial responses and branches $w$ parallel blocks from each one [27, 5, 41, 6, 37]. Given a prompt $x$, beam search first generates $N$ parallel blocks. However, unlike BFS, beam search keeps the top $B$ beams with the highest scores, and then samples $w$ parallel blocks per beam at the next step. Since $B\times w=N$ blocks are sampled at each step, the compute budget is also $N$. We illustrate beam search with $N=4$ and $w=2$ in Fig.˜2 (right).
@@
We remark a crucial detail of beam search and DVTS is how the final set of beams/responses are aggregated. Prior works [6, 37, 35] select the response with the highest score, which is analogous to a final best-of-$N$ (BoN; Algorithm˜2). Instead, we found that taking a weighted majority vote (WMV; Algorithm˜3) led to much better performance, which is demonstrated by Fig.˜3 (left).
Computational Efficiency of Block-wise Search. Since value scores are only used at the end of each block or the end of the whole response, the FLOPs required for block-wise value model guidance is a tiny fraction ($\ll 1\%$) of the generation cost from $\pi$. We compute FLOP estimates in Appendix˜H to concretely show this.
@@ 3 Experiments
We extensively evaluate value-guided search (VGS) with our 1.5B value model DeepSeek-VM-1.5B, focusing on guiding the CoT reasoning of DeepSeek models [19]. The best VGS setup for our value model is beam search with final WMV aggregation, beam width 2, block size 4096 and with DVTS (for larger inference budgets). We show this setup outperforms other test-time compute methods (e.g., MV, WMV, BoN) and other scoring models (e.g., existing 7B PRMs and a 1.5B Bradley-Terry reward model trained on our dataset). We remark our search results use a fixed beam width and block size for all problems; this is more practical than prior works on “compute-optimal scaling” which vary search parameters for each problem and require estimating each problem’s difficulty [6, 37, 26]. Please see Appendices˜E and F for additional details on value model training and inference.
Benchmarks. We evaluate on the American Invitational Mathematics Exam (AIME) and the February Harvard-MIT Mathematics Tournament (HMMT). [^2] https://maa.org/maa-invitational-competitions and https://www.hmmt.org Both AIME and HMMT are prestigious high school math competitions in the US that have also been used to evaluate frontier LLMs [32, 19, 1]. We use AIME I & II and the individual part of HMMT, yielding $30$ problems per competition. We selected the block size and beam width based on the performance AIME-24 as the validation set (ablations in Section˜4.1). Then, we evaluate on AIME-25 and HMMT-25 since they happened after the release of DeepSeek and OpenR1. This controls for data contamination since the test problems were released after the underlying reference models and datasets. We also ensure that there are no problems with $50\%$ 8-gram overlap between training and test sets as a further sanity check for contamination. Unless otherwise stated, we report the overall accuracy averaged across AIME-25 and HMMT-25. In Appendix˜C, we report individual per-benchmark results for the 2024 and 2025 editions of both math competitions.
Baseline Models. We evaluate two state-of-the-art 7B PRMs with distinct training styles: Math-Shepherd-Mistral-7B-PRM [45] and Qwen2.5-Math-PRM-7B [51]. Math-Shepherd uses Monte-Carlo roll-outs from each step to estimate per-step value while the Qwen2.5 PRM uses LLM-Judge annotation for each step, similar to the per-step human annotation of PRM800K [24]. As a step-level value model, Math-Shepherd-PRM-7B is more related to our token-level value model. Finally, we also evaluate a 1.5B Bradley-Terry (BT) [8] model, called DeepSeek-BT-1.5B, which we trained using our dataset (see Appendix˜G for training details).
Test-time scaling DeepSeek-1.5B ($N=256$) AIME-25 HMMT-25 AVG
VGS w/ DeepSeek-VM-1.5B (ours) $46.7\pm 0.7$ $32.8\pm 0.8$ $\mathbf{39.8\pm 0.5}$
WMV w/ DeepSeek-VM-1.5B (ours) $45.1\pm 2.2$ $28.9\pm 2.6$ $37.0\pm 1.7$
VGS w/ DeepSeek-BT-1.5B (ours) $40.6\pm 0.8$ $27.5\pm 0.0$ $34.1\pm 0.4$
WMV w/ DeepSeek-BT-1.5B (ours) $40.5\pm 2.9$ $24.6\pm 4.7$ $32.6\pm 1.6$
VGS w/ Qwen2.5-Math-PRM-7B $38.9\pm 1.4$ $24.2\pm 0.2$ $31.6\pm 0.7$
WMV w/ Qwen2.5-Math-PRM-7B $39.1\pm 2.1$ $24.0\pm 3.2$ $31.6\pm 1.9$
VGS w/ MathShepherd-PRM-7B $41.9\pm 1.4$ $23.9\pm 1.4$ $32.9\pm 1.0$
WMV w/ MathShepherd-PRM-7B $40.0\pm 2.5$ $25.6\pm 3.1$ $32.8\pm 2.0$
MV@256 $38.9\pm 1.9$ $24.3\pm 2.9$ $31.6\pm 1.7$
Test-time scaling DeepSeek-7B ($N=128$)
VGS w/ DeepSeek-VM-1.5B $59.4\pm 0.8$ $41.1\pm 1.6$ $\mathbf{50.3\pm 0.9}$
MV $56.5\pm 1.6$ $33.8\pm 2.5$ $45.2\pm 1.5$
@@
DeepSeek-1.5B Pass@1 $22.4\pm 4.1$ $13.0\pm 3.9$ $17.7\pm 2.8$
DeepSeek-32B Pass@1 $60.4\pm 6.0$ $42.1\pm 5.2$ $51.3\pm 4.0$
Deepseek-R1 (671B) Pass@1 $70.0\pm 0.9$ $46.7\pm 2.4$ $\mathbf{58.4\pm 1.3}$
Table 1: (Top) Weighted majority vote (WMV) and VGS results for DeepSeek-1.5B with an inference budget of $N=256$, using various scoring models. (Middle) Compares MV and VGS for larger DeepSeek models guided with our DeepSeek-VM-1.5B. (Bottom) Lists performance of DeepSeek models and strong close-sourced reasoning models. For VGS, $\pm$ indicates standard deviation across 3 seeds; for MV, WMV, Pass@N, $\pm$ denotes bootstrap with $100$ repetitions. We bold the highest avg. accuracy and underline second highest. Section˜C.1 contains more baselines.
3.1 Main Results (Table˜1)
In the top section of Table˜1, we fix the generator to DeepSeek-1.5B [^3] Throughout the paper, we use DeepSeek-XB as shorthand for DeepSeek-R1-Distill-Qwen-XB. and test-time budget to $N=256$, and compare VGS to weighted majority voting (WMV), using our value model, the BT model and baseline PRMs. We see that VGS and WMV with DeepSeek-VM-1.5B achieve the two highest scores, outperforming the BT reward model and prior PRMs. This shows that our value model is not only a strong outcome reward model (ORM) but also an effective value model for guiding search. Intriguingly, while DeepSeek-BT-1.5B was only trained as an ORM, we find that VGS also improves performance relative to WMV, suggesting that BT models may also provide meaningful block-wise feedback to guide search. We also observe that accuracies for the 7B baseline PRMs (MathSheperd and Qwen2.5-Math) are only slightly higher than MV@256 and do not improve with search, which suggests that these PRMs are likely out-of-distribution (OOD) for the long CoTs generated by DeepSeek-1.5B.
In the middle section of Table˜1, we guide the stronger 7B DeepSeek model and compare VGS to MV, a standard TTC method that does not use an external scoring model. We see that VGS again achieves higher accuracy than MV for 7B, which suggests that DeepSeek-VM-1.5B is also useful for guiding CoT of larger DeepSeek models. In Table˜2, we also include results for guiding 14B DeepSeek, where we observe that the gap between VGS and MV becomes much smaller. This suggests that DeepSeek-14B CoTs may be becoming OOD for our value model trained on DeepSeek-1.5B CoTs. To guide more capable models, new value models should be trained on rollouts from similarly capable models; we however do not foresee this being a practical concern given the scalability of our training process (described in Section˜2 and summarized in Appendix˜B).